Juanita, who is 1.82 meters tall, wants to find the height of a tree in her backyard. From the tree's base, she walks 12.2 meters along the tree's shadow until her head is in a position where the tip of her shadow exactly overlaps the end of the treetop's shadow. She is now 6.1 meters from the end of the shadows. How tall is the tree?
Please help, thank you.

Answer by checkley77(12844)

1.82/6.1=x/(12.2+6.1)
1.82/6.1=x/18.3 cross multiply
6.1x=1.82*18.3
6.1x=33.306
x=33.306/6.1
x=5.46 meters is the height of the tree.